OpenStudy (anonymous):

If 1.0 gram of matter was converted to energy, how much energy would be formed? 9.0 x 10-13 joules 9.0 x 1013 joules 3.0 x 1013 joules 3.0 x 108 joules

OpenStudy (anonymous):

@Nnesha @welshfella @Mehek14 @JoannaBlackwelder @DullJackel09 @ganeshie8 PLEaseee

OpenStudy (anonymous):

@JoannaBlackwelder Please help me

OpenStudy (anonymous):

i know you use e=mc^2

OpenStudy (anonymous):

but I'm not sure what everything stands for

OpenStudy (frostbite):

E is the energy c is the speed of light \(c\approx3 \times 10^{8}~m/s\) m is the mass. In order to obtain the answer in joule you need to convert gram to kilogram as: \[\large 1~J=1~\frac{ kg \times m^2 }{ s^2 }\]

OpenStudy (frostbite):

I help you til you get the question correct. All you got to do really, is to convert your 1 gram to kilogram and then plug into the equation using the speed of light I gave you.

OpenStudy (anonymous):

Thank you SO much!!! this helped a lot

OpenStudy (frostbite):

Well no problem at all!

OpenStudy (frostbite):

Found the correct answer? :)